### v1

# Quantum thermalization and the route to ergodicity

Abstract
We consider a minimal model for quantum thermalization of coupled chaotic subsystems. The route towards ergodicity is explored as a function of the coupling strength. The results are contrasted with the predictions of standard Random Matrix Theory (RMT) and the Eigenstates Thermalization Hypothesis (ETH). We highlight a coupling regime of disparity between the spectral statistics that indicates chaos, and ergodicity measures that indicate lack of ETH thermalization. The analysis involves a revision of the energy shell concept, in a way that is consistent but independent of the semiclassical perspective.
Random matrix theory (RMT), and specifically the theory of Wigner banded random matrices (WBRM) [1, 2], have been proposed as a framework to describe ergodization and thermalization. Consider the quantum thermalization between coupled subsystems of a bipartite disordered system, where each subsystem by itself is generically chaotic. The Hamiltonian is ${H=H_{0}+\lambda V}$ where the inter-system interaction $\lambda V$ mixes the eigenstates of the uncoupled subsystems. It has been previously realized that generic perturbations of a quantized chaotic system (QCS) can be modeled by WBRM [3, 4, 5, 6, 7, 8, 9]. It follows that for relatively weak coupling the mixing of the levels is described by the Wigner Lorentzian, while for strong coupling the line-shape becomes non-perturbative. Namely, for WBRM it is a semicircle line-shape, while for a QCS it is a semiclassical line-shape that reflects the underlying phase-space geometry. The latter is implied by the association of chaotic eigenstates with micro-canonical distributions in Berry’s conjecture and the eigenstate thermalization hypotheis (ETH) [10, 11, 12, 13, 14, 15, 16].
As shown below, the application of WBRM theory to the thermalization of two chaotic subsystems is rather subtle. For illustration we consider a bi-partite Bose-Hubbard model (BHM) [17, 18, 19, 20, 21, 22, 23, 24]. The eigenstates of the disconnected subsystems $H_{0}|x,n\rangle=\varepsilon_{x,n}|x,n\rangle$ are direct products of subsystem eigenstates with well-defined integer subsystem occupations, indexed by two good quantum numbers: the occupation difference $x$ and the index $n$ of the state within the factorized spectrum (for a given $x$). These are plotted in Fig. 1, where each point represents an unperturbed eigenstate. The coupling $\left\langle x,n\middle|V\middle|x^{\prime},n^{\prime}\right\rangle$ connects only states with ${|x-x^{\prime}|=1}$. This 2D geometry of the basis for the BHM is very different from the 1D geometry of the WBRM (left panel of Fig. 1), with substantial implications on the route to ergodicity.
Figure 1: The unperturbed basis. The unperturbed energies $\varepsilon_{x}$ and $\varepsilon_{x,n}$ for the WBRM (a) and for the BHM (b,c), respectively. The bandwidth $\Delta_{b}$ is the maximal range in energy that can be reached by a single transition from an initial state. Possible first-order transitions are indicated. The width of the energy-shell $\Delta x_{\text{shl}}$ is defined via the microcanonical distribution $p_{x}$ (d). It is not sensitive to $\lambda$ for the BHM, and has a well defined finite value in the limit of zero coupling. In contrast, the WBRM model can be regarded as a singular limit that features ${\Delta x_{\text{shl}}\propto\lambda}$.
Outline.– We introduce a critical discussion of the energy-shell concept. Then we explore the route towards ergodicity for the BHM and contrast it with the WBRM.
The BHM.– Consider the bipartite BHM for $N$ particles in $M=M_{A}{+}M_{B}$ sites, where $M_{\alpha}$ is the number of sites in subsystem $\alpha=A,B$. The Hamiltonian is,
$H=\sum_{\alpha=A,B}H_{\alpha}+\lambda V$ (1)
where the subsystem Hamiltonians are,
$H_{\alpha}$ $=$ $\sum_{i}\varepsilon_{\alpha,i}\hat{n}_{\alpha,i}+\frac{U}{2}\sum_{i}{\hat{n}}_{\alpha,i}({\hat{n}}_{\alpha,i}{-}1)$
$-K\left(\sum_{\langle i,j\rangle}{\hat{a}}_{\alpha,i}^{\dagger}{\hat{a}}_{\alpha,j}+{\rm H.c.}\right)$ (2)
and the inter-system coupling is,
$V={\hat{a}}_{A,1}^{\dagger}{\hat{a}}_{B,1}+{\hat{a}}_{A,M_{A}}^{\dagger}{\hat{a}}_{B,M_{B}}+{\rm H.c.}$
In the above, the boson operator ${\hat{a}}_{\alpha,i}$ annihilates a particle in the $i$-th site of the $\alpha$ subsystem with population ${\hat{n}}_{\alpha,i}={\hat{a}}_{\alpha,i}^{\dagger}{\hat{a}}_{\alpha,i}$, the intra-system hopping frequency is $K$, and $\langle i,j\rangle$ denotes nearest-neighbor pairs. The coupling strength is $U$, and the weak on-site random potential $\varepsilon_{\alpha,i}$ breaks parity. In the numerical demonstrations we have ${M_{A}{=}M_{B}{=}4}$, $N{=}10$, and ${|\varepsilon_{\alpha,i}|<0.05}$, while ${K{=}1}$, and $U{=}1$.
The Hilbert-space dimension of the BHM is,
$D_{M}(N)=\frac{(N{+}M{-}1)!}{N!(M{-}1)!}=\sum_{x=-N}^{N}D_{x}$ (3)
where $D_{x}=D_{M_{A}}\left(\frac{N{+}x}{2}\right)\times D_{M_{B}}\left(\frac{N{-}x}{2}\right)$ is the dimension of the $x$ sector of the spectrum (i.e. the number of states in each column of the diagram in Fig. 1) and $x=-N,-N{+}2,\cdots,N{-}2,N$. In the absence of inter-system coupling ($\lambda=0$), the eigenenergies of the unperturbed Hamiltonian $H_{0}=H_{A}+H_{B}$ are $\varepsilon_{x,n}$, where the running index ${n=1,\dots,{{\color[rgb]{0.6,0,0.1}D_{x}}}}$ labels the $x$-sorted unperturbed energies.
Eigenstates of the coupled subsystems.– Once the coupling $\lambda$ it turned on, the subsystems can exchange particles, and the coordinate $x$ is no longer a good quantum number. Accordingly we use a running index ${\nu=1,\cdots,D_{M}(N)}$ to label the many-body eigenstates and the respective eigenenergies $E_{\nu}$. In the presentated results below, these energies are scaled as ${E:=(E{-}E_{\rm min})/(E_{\rm max}{-}E_{\rm min})}$, where $E_{\rm min}$ and $E_{\rm max}$ are the ground and top energies.
Each exact eigenstate features a probability distribution over the unperturbed states, namely,
$p_{x,n}^{(\nu)}\ \ =\ \ \Big{|}\left\langle x,n\middle|E_{\nu}\right\rangle\Big{|}^{2}$ (4)
The associated population-imbalance and energy distributions are,
$p_{x}^{(\nu)}$ $=$ $\ \sum_{n}p_{x,n}^{(\nu)}$ (5)
$p^{(\nu)}(\varepsilon)$ $=$ $\ \sum_{x,n}p_{x,n}^{(\nu)}\,\delta(\varepsilon-\varepsilon_{x,n})$ (6)
Using the above probability distributions, we define the following measures for the eigenstates, characterizing their spreading over the unperturbed basis:
$P_{\nu}$ $=$ $\max_{x,\nu}\Big{\{}p_{x,n}^{(\nu)}\Big{\}}\ =\ \text{Peak probability}$ (7)
$\mathcal{M}_{\nu}$ $=$ $\left[\sum_{x,n}\big{|}p_{x,n}^{(\nu)}\big{|}^{2}\right]^{-1}\!\!=\text{Participation number}$ (8)
$\Delta x_{\nu}$ $=$ dispersion of $x$ via Eq. (5) (9)
$\Delta{\varepsilon}_{\nu}$ $=$ dispersion of $\varepsilon$ via Eq. (6) (10)
The peak probability indicates whether the exact eigenstate $|E_{\nu}\rangle$ is a first-order perturbed eigenstate $|x,n\rangle$ of the uncoupled system. The participation number estimates the number of unperturbed eigenstates that contribute to the exact eigenstate. The dispersions quantify the spread in $x$ and in $\varepsilon$. It should be noted that while the full energy dispersion of an exact eigenstate is identically zero, $\Delta\varepsilon_{\nu}$ is the dispersion over the unperturbed energies, i.e. $\Delta\varepsilon_{\nu}=[\langle H_{0}^{2}\rangle_{\nu}-\langle H_{0}\rangle_{\nu}^{2}]^{1/2}$
Figure 2: The width of the energy shell. (a) The width $\Delta x_{\text{shl}}$ versus the width $\delta E$ of the arbitrary energy window about ${E=0.41666}$ for $\lambda=10^{-3}$ (solid blue line), $10^{-2}$ (dashed red), $10^{-1}$ (dotted orange), and 1.01.01.0 (dash-dotted purple). The width of the energy shell is well defined (assuming quantum mechanically large but classically small $\delta E$) and the dependence of the obtained shell width on the coupling strength $\lambda$ is weak. (b,c) The dependence of $\Delta x_{\rm shl}$, ${\cal M}_{\rm shl}$ and $\Delta\varepsilon_{\rm shl}$ on the coupling $\lambda$. In the latter case one observe a residual value for small $\lambda$, corresponding to the arbitrary $\delta E$ (dashed line is for $\delta E=0.0333$ whereas dash-dotted line is for $\delta E=0.0166$). This numerical artifact can be eliminated by looking on ${\overline{\Delta\varepsilon_{\nu}}}$ (solid line), even if the eigenstates are non-ergodic! The dotted line is a guide to the eye, corresponding to linear $\Delta\varepsilon_{\rm shl}(\lambda)$ dependence
The energy shell.– Semiclassically, a microcanonical distribution occupies an energy shell of thickness $\delta E$ in phase-space. In the classical context the ‘energy surface’ limit $\delta E\rightarrow 0$, is well defined. Quantum mechanically, the thickness $\delta E$ should be large with respect to the mean level-spacing but small with respect to the energy scale on which this spacing varies, i.e. the shell width should be ‘classically small but quantum mechanically large’.
It is practically useful to project the shell either over $x$ or over $\varepsilon$ to obtain the shell probability distributions $p_{x}$ and $p(\varepsilon)$. This means that $p_{x}\equiv\overline{p^{(\nu)}_{x}}$ and $p(\varepsilon)\equiv\overline{p^{(\nu)}(\varepsilon)}$ are obtained by respectively averaging $p_{x}^{(\nu)}$ and $p^{(\nu)}(\varepsilon)$ over eigenstates $\nu$ that belong to the $\delta E$ energy window, i.e. ${|E_{\nu}-E|<\delta E/2}$. We define the dispersion $\Delta x_{\rm shl}$ as the second moment of the distribution $p_{x}$, as illustrated in Fig. 1. Similarly, we extract the dispersion $\Delta\varepsilon_{\rm shl}$ from the unperturbed energy distribution $p(\varepsilon)$.
In Fig. 2, we plot the dependence of $\Delta x_{\rm shl}$ and $\Delta\varepsilon_{\rm shl}$ on the arbitrary chosen window $\delta E$ and the coupling strength $\lambda$. When $\delta E$ is large with respect to the level spacing, the shape of $p(x)$ becomes independent of $\delta E$, and only very weakly dependent on $\lambda$. Thus, the dispersion $\Delta x_{\rm shl}$ is well defined.
By contrast, the energy dispersion $\Delta\varepsilon_{\rm shl}$ approaches the arbitrary window width $\delta E$ at small $\lambda$. This numerical artifact indicating that $\delta E$ is still too large. The dispersion $\Delta\varepsilon_{\rm shl}$ is the convolution of the individual widths $\Delta\varepsilon_{\nu}$ with the arbitrary $\delta E$. Physically meaningful result is obtained provided the contribution of $\delta E$ is negligible. It turns out that this numerical limitation can be avoided, because for classically small $\lambda$ we get [27]
$\Delta\varepsilon_{\rm shl}=\overline{\Delta\varepsilon_{\nu}}=\lambda\sqrt{[V^{2}]_{\varepsilon,\varepsilon}-[V]_{\varepsilon,\varepsilon}^{2}}$ (11)
This holds even if the eigenstates are non-ergodic, as further discussed below under ‘False ergodicity measures’. Fig. 2c demonstrates the validity of this statement, and the implied linear dependence of $\Delta\varepsilon_{\rm shl}$ on $\lambda$.
Distinguishing ergodicity from chaos.– It is a common misconception to identify “quantum chaos” with ”quantum ergodicity”. Quantum chaos is identified by the spectral statistics. In Fig. 3 we display a re-scaled version of a common measure for level repulsion, namely the mean ratio of consecutive level spacings [25],
$\tilde{r}\equiv\frac{\langle r\rangle-r_{\rm Poisson}}{r_{\rm GOE}-r_{\rm Poisson}}$ (12)
The value $\langle r\rangle=r_{\rm Poisson}=0.38629$ (${\tilde{r}}{=}0$) characterizes spectra that consist of uncorrelated levels. This applies to integrable systems but also to the uncoupled chaotic systems of our model. The value $\langle r\rangle=r_{\rm GOE}=0.53590$ (${\tilde{r}}{=}1$) holds for spectra of fully chaotic systems with Wigner-Dyson level spacing statistics.
The proper measure for the quantum ergodicity of the eigenstates, is the mean participation number, $\overline{\mathcal{M}_{\nu}}/\mathcal{M}_{\rm shl}$ (dashed line in Fig. 3). The $\mathcal{M}_{\nu}$ are calculated for the individual eigenstates, while $\mathcal{M}_{\rm shl}$ is calculated from the $p_{x,n}$ of the energy shell. For GOE statistics, a value of $\overline{\mathcal{M}_{\nu}}/\mathcal{M}_{\rm shl}\approx 1/3$ indicates the ergodicity of the eigenstates. The 1/31/31 / 3 factor emerges because the $p_{x,n}^{(\nu)}$ probabilities are not uniformly spread over the energy shell but statistically fluctuate.
Comparing the two measures, we note that the ergodic regime in Fig. 3 is much smaller than the chaotic regime indicated by spectral statistics. For a substantial $\lambda$ range the system is chaotic, yet its eigenstates are not ergodic.
The ETH ergodicity measures.– The ETH predicts that for ergodic eigenstates $|E_{\nu}\rangle$ the expectation value of any local observable, such as $\left\langle x\right\rangle_{\nu}$, will exhibit low state-to-state fluctuations, with dispersion ${\sigma_{x}\propto 1/\sqrt{\mathcal{M}}}$. The mere decrease of state-to-state fluctuations however, is not a sufficient indication for ergodicity, as clearly demonstrated by comparing the earlier (weak coupling) drop of $\sigma_{x}$ with the later (stronger coupling) rise of $\overline{{\cal M}_{\nu}}$ in Fig. 3.
Figure 3: Chaos and Ergodicity Measures. Misc measures are calculated for the energy shell ${0.4<E<0.4333}$ as a function of the coupling strength $\lambda$, and normalized per the expected GOE values. The spectral measure for chaos $\tilde{r}$ (square markers) is correlated with $\overline{\Delta x_{\nu}}/\Delta x_{\rm shl}$ (solid line). The ETH measure $\sigma_{x}/\Delta x_{\rm shl}$ (dotted line) is an insufficient measure for ergodization. The measure $\overline{\mathcal{M}_{\nu}}/(\mathcal{M}_{\rm shl}/3)$ (dashed line) is an effective measure for the actual identification of the ergodic regime. We also plot the mean peak probability $\overline{P_{\nu}}$ (dash-dotted line), which is used to identify the first-order perturbative regime (arbitrarily defined as $\overline{P_{\nu}}>0.6$). Inverted triangles mark the coupling strengths for the states presented Fig. 4.
[…]
Contrast with Wigner theory.– Referring to the 2D energy landscape of the BHM, see Fig. 1, the width $\Delta x_{\rm shl}$ is finite, with negligible dependence on $\lambda$, see Fig. 2. This should be contrasted with the WBRM paradigm, where the horizonatal and the vertical dimensions are degenerate. Namely, $\Delta x_{\rm shl}=\Delta\varepsilon_{\rm shl}/\Delta_{0}$, where $\Delta_{0}$ is the mean level spacing. Thus, in the WBRM model, $x$ can be viewed as a dimensionless energy coordinate. This observation leads to the understanding that the mathematical-oriented WBRM model is an oversimplification. The actual route towards thermalization in a realistic physical system like the BHM is more complicated and non-universal.
The route to ergodicity.– We identify five distinct physical regimes in the intricate progression of the BHM towards ergodicity, as the coupling strength $\lambda$ increases. These are marked by the shaded regions in Fig. 3, with a representative eigenstate for each in Fig. 4. For the general discussion we note that the relevant model parameters are the mean level spacing $\Delta_{0}$, the bandwidth $\Delta_{b}$, and the typical value $\bar{V}$ of the in-band matrix elements. It is useful to define dimensionless bandwidth ${b=\Delta_{E}/\Delta_{0}}$, and dimensionless coupling ${\tilde{\lambda}:=\lambda\bar{V}/\Delta_{0}}$.
(a) The perturbative regime.– For very small values of $\lambda$, The peak probability $P_{\nu}$ is of order unity, and the $|E_{\nu}\rangle$ eigenstates can be described by 1st-order perturbation theory, which implies
$p_{x,n}^{(\nu)}=\left|\frac{\left\langle x,n\middle|\lambda V\middle|x_{0},n_{0}\right\rangle}{\varepsilon_{x,n}-\varepsilon_{x_{0},n_{0}}}\right|^{2},\ \ \ (x,n)\neq(x_{0},n_{0})$ (13)
where $(x_{0},n_{0})$ indicate the unperturbed eigenstate. This approximation holds as long as ${\tilde{\lambda}<1}$. It has been argued in [15] that Eq. (13) holds for the tails of the coarse-grained $p_{x,n}^{(\nu)}$ for larger values of $\lambda$, as long as quantum ergodicity is not yet attained. The validity of this claim for the BHM is demonstrated with striking accuracy in the lower panels of Fig. 4.
(b) The Mixing regime.– As $\lambda$ is increased nearby levels are mixed, leading to level repulsion reflected by the spectral measure $r$. At this stage the eigenstates spread in the ’horizontal’ $x$ direction of the ${(x,\varepsilon)}$ energy shell, yet remain localized in the ’vertical’ $\varepsilon$ direction. A rough estimate for this horizontal mixing is provided by the standard Anderson model, predicting a localization length ${\xi\sim\lambda^{2}}$. Full horizontal mixing is attained once ${\xi\sim\sin N}$. This implies a border ${\tilde{\lambda}_{\rm mix}\sim\sqrt{N}}$ at which ’horizontal ergodicity’ $\overline{\Delta x_{\nu}}/\Delta x_{\rm shl}\approx 1$ is attained.
(c) The Wigner regime.– Once the horizontal mixing is achieved, we get into the Wigner regime of vertical spreading, where the WBRM theory applies. A terse summary of the WBRM theory is provided in the [27]. The mixed levels form a core of vertical width ${\Gamma_{E}\propto\lambda^{2}}$. This core determines the participation number $\mathcal{M}_{\nu}$. Full ergodicity is reached once ${\Gamma_{E}}$ becomes comparable with the vertical width of the energy shell Eq. (11). Consequently, the expected ergodicity threshold is ${\tilde{\lambda}_{\rm erg}\sim\sqrt{b}}$.
(d) The Ergodic regime.– For $\lambda>\lambda_{\rm erg}$ there is full mixing within the energy shell, both horizontally and vertically. This regime is semiclassical. To the extent it conforms to the WBRM model, we can say that in the Wigner regime ${\Gamma_{E}\ll\Delta_{E}<\Delta_{b}}$, while in the ergodic regime ${\Gamma_{E}\sim\Delta_{E}>\Delta_{b}}$, where $\Gamma_{E}$ indicates the width of the ‘core’ region of non-perturbative mixing.
(e) The Integrable regime.– Finally, when ${\lambda\gg K}$ the intra-system hopping may be neglected and the system is reduced to the sum of disconnected integrable Bose-Hubbard dimers. Consequently, ergodicity is lost, $M_{\nu}$ and $\Delta\varepsilon_{\nu}$ decrease, and the state-to-state variance $\sigma_{x}$ increases as the ETH is violated. This regime is non-universal. In the WBRM the somewhat analogous regime is the Anderson regime ${\tilde{\lambda}>b^{3/2}}$, see [27] for details.
False ergodicity measures.– Unlike $\overline{\mathcal{M}_{\nu}}/\mathcal{M}_{\rm shl}$, the ratio $\overline{\Delta\varepsilon_{\nu}}/\Delta\varepsilon_{\rm shl}$ can not be used as an ergodicity measure. In fact, the first equality in Eq. (11) implies that the latter ratio is of order unity, irrespective of ergodicity. A rather straightforward argument (see [27]) shows that ${\Delta\varepsilon_{\nu}}$ for any eigenstate is determined by the strength of the perturbation $V$, and is linear in $\lambda$. This is true even if the eigenstate is localized within the energy shell. The most dramatic demonstration is based on Eq. (13) for which ${\mathcal{M}\sim 1}$, but nevertheless the dispersion ${\Delta\varepsilon_{\nu}}$ is determined by the tails, and does not care about ${\mathcal{M}}$. Consequently, the linear $\lambda$ dependence is not affected by ${\tilde{\lambda}_{\rm mix}}$ neither by the ${\tilde{\lambda}_{\rm erg}}$ threshold, and cannot be used as an ergodicity measure.
The regime borders.– Consider the scaling of the mixing- and ergodicity borders $\lambda_{\rm mix}$ and the $\lambda_{\rm erg}$, with $N$. Within a semiclassical treatment a rough estimate for the localization length is ${\xi=(1/\Delta_{0})D_{x}}$, where the diffusion coefficient in $x$ is ${D_{x}=N^{2}D_{cl}(\lambda)}$. Note that $D_{cl}\propto[\lambda V_{cl}]^{2}$, where $V_{cl}$ is the strength of perturbation in ‘classical’ units. The density of states is ${(1/\Delta_{0})=N^{d}\Omega_{cl}}$, where the number of degrees of freedoms is $d=M{-}1$, and $\Omega_{cl}$ is the phase-space area of the energy shell. We note that $\xi$ is possibly larger as discussed in [23]. From ${\xi<N}$ we get the condition
$\Omega_{cl}D_{cl}(\lambda)<(1/N)^{1+d}\ \ \ \ \leadsto\ \ \ \ \lambda_{\rm mix}$ (14)
Turning to the ergodic border, the condition is ${\Gamma_{E}<\Delta_{b}}$. The Feingold-Peres relation [3, 4, 5, 6] (equivalent to ETH) is Eq.6 of [15] that relates matrix elements to fluctuations. It is implied that $\Delta_{b}=\hbar/\tau_{cl}$, and that ${\Gamma_{E}=(1/\Delta_{0})\lambda^{2}\bar{V}^{2}=\lambda^{2}V_{cl}^{2}\tau_{cl}/\hbar}$. Here $\hbar=1/N$. Thus we get the condition
$\lambda^{2}V_{cl}^{2}\tau_{cl}^{2}<(1/N)^{2}\ \ \ \ \leadsto\ \ \ \ \lambda_{\rm erg}$ (15)
Comparing the two conditions we conclude that for large $N$ the mixing border is typically encountered before the ergodic border.
Summary.– The thermalization of large disordered system can be modeled as the ergodization of small nearby chaotic subsystems, sometimes referred to as hot spots [26]. RMT modeling provides a partial understanding of the route towards ergodization. For physical subsytems a more careful definition of the energy-shell concept is required. Then it becomes clear that two distinct steps are involved, with different thresholds $\lambda_{\rm mix}$ and $\lambda_{\rm erg}$, that signify spectral quantum chaos and ETH ergodization respectively.
Acknowledgments – AV acknowledges support from the NSF through a grant for ITAMP at Harvard University. DC acknowledges support by the Israel Science Foundation, grant No.518/22.